# Cacophony: An Improved Contrastive Audio-Text Model

### v2

@@ III Neural Architecture
Figure 3: The proposed system consists of a two-stage process, as illustrated in the system diagram: Left (Stage 1): Audio-MAE Pre-pretraining: This stage is conducted on our collected dataset. After this stage, the audio decoder is discarded, and the encoder is retained and initialized as the audio encoder for the second stage of training. Right (Stage 2): Contrastive-Captioning Objective Pretraining: This stage involves training three components – the audio encoder, the text encoder, and the text decoder. The second stage is dedicated to achieving a contrastive-captioning objective, aligning and fine-tuning the interaction between the audio and text components.
We train Cacophony in two stages, as depicted in Fig. 3. First, we train an audio encoder with the MAE objective, using audio-only data from our collected large-scale dataset. Second, we take this trained audio encoder and use it as the initilzation for training the audio-text model with both contrastive and captioning objectives. The first audio-only MAE stage is also called pre-pretraining [36], while we refer to the second audio-text contrastive and captioning stage simply as pretraining.
III-A Stage 1: Audio encoder pre-pretraining
We pre-pretrain our audio encoder with an MAE objective, which involves masking random patches of the input signals (e.g. images or spectrograms) and then reconstructing these masked portions [23, 24]. Specifically, in the MAE training, a transformer encoder initially processes the unmasked patches. Following this, a transformer decoder, which will be discarded before the second-stage contrastive-captioning training, predicts both the masked and unmasked regions. This MAE objective encourages the learning of global, contextualized representations over arbitrary subsets of spectrogram patches [24]. Compared to methods based on supervised classification objectives, MAE does not required labeled data. It also demonstrates improved performance with increased model size [23] and increased dataset size [36].
@@ III-A Stage 1: Audio encoder pre-pretraining
III-B Stage 2: Contrastive-captioning pretraining
Given the audio encoder trained in the first stage, we now introduce a text encoder/decoder and train a contrastive model with an auxiliary captioining objective, in a setup similar to BLIP [26] and CoCa [25]. For the contrastive objective [39, 12], we use learned linear projections to map the text and audio embeddings to the same dimension, followed by an $l_{2}$-normalization layer. Then, we use matched audio-text embeddings as positives and all non-paired examples as hard negatives for the pairwise InfoNCE loss as described in [40]:
@@ III-B Stage 2: Contrastive-captioning pretraining
where $x_{i}$ and $y_{j}$ are $l_{2}$-normalized embeddings of the audio in the $i$-th pair and text in the $j$-th pair respectively and $\tau$ is a learnable temperature.
@@
where here $y_{t}$ is the $t$-th ground-truth token for a given caption $y$ and $T$ is the caption’s total length. As a result, we apply both contrastive and generative objectives in the second stage training as follows:
@@
where $\lambda_{Cap}$ is hyper-parameter representing relative weight.
Architecturally, the second-stage training involves training three key modules: an audio encoder, text encoder, and text decoder. For the audio encoder, we use the encoder from the first stage of AudioMAE pre-pretraining, outlined in Sec. III-A.
Our text encoder is a transformer that follows RoBERTa’s architecture. However, unlike RoBERTa, we apply causal self-attention rather than bidirectional self-attention. This is a necessary modification to prevent information leakage to the text decoder. We initialize our model with the RoBERTa pretrained weights. The outputs from audio encoder and text encoder are framewise and token-wise embeddings respectively; in order to obtain a single embedding vector for each modality, we choose to integrate multi-head attention poolers on top of the sequential embeddings [41, 25] for the contrastive objective, shown in the right of Fig. 3.
Our text decoder consists of stacked transformer layers on top of our text encoder, again with causal self-attention to prevent next-token prediction leakage. Each transformer block has a cross-attention layer which attends to the full output of our audio encoder. The text decoder layers are initialized randomly.
@@
IV Training Details
IV-A Training setup
As outlined in Sec. II, the training datasets are categorized into three types: the ‘clean-labeled’ dataset consists of 1,212 hours, the ‘noisy labeled’ dataset consists of 3,003 hours, and the ‘weakly/unlabeled’ dataset consists of 9,017 hours. All audio files were processed into a mono channel with a sampling rate of 16 kHz. We extract mel-spectrograms using a 25 ms window and a 10 ms hop length, with 128 mel bands and 512 as FFT size. The spectrograms are then converted into 16$\times$16 patches without overlap.
In the first-stage training, we randomly sample 15 seconds of audio for each recording, corresponding to a patch sequence length of 750. At the second stage of training, we limit the audio length to 10.24 seconds, equivalent to a patch sequence length of 512, i.e. when the patch sequence length exceeds 512, we randomly sample 512 patches from the full sequence, following PaSST [42]. Text samples are tokenized with the pretrained Roberta tokenizer and truncated to maximum length 77. Our audio encoder, employed in both training stages, is a 12-layer ViT-Base (ViT-B) Transformer with 8 attention heads, a hidden size of 768, an intermediate size of 3072, and utilizes the SiLU activation function. Additionally, a 12-layer decoder with the same settings is used during the MAE training phase. For the attention pooler, we choose to use $n_{head}=8$. Our codebase is developed using the JAX framework [43] with the Flax library  [44] and the model is trained on a TPU-v3 Pod in both stages. In particular, we employ data parallelism to shard both the text and audio data along all the individual TPU cores, process the audio and text encoding in parallel, and then all-gather the pooled embeddings in order to compute the contrastive similarity matrix (though the matrix multiplies themselves are sharded over all devices).
When training AudioMAE, we use a batch size of 512, a learning rate of $2\times 10^{-4}$, and employ the AdamW [45] optimizer with a weight decay of 0.01. We train for 200,000 steps, with a 10,000 step learning rate warm-up and a cosine decay to $1\times 10^{-6}$. In the contrastive stage, we use a batch size of 4096. We again use AdamW with weight decay 0.01. The learning rate is warmed up for 10,000 to a peak of $1\times 10^{-5}$ and is decayed to $1\times 10^{-6}$ over 300,000 (total) steps following a cosine schedule.
@@ IV-B Sharpness-aware minimization
Given our use of a large training batch size of 4096 for the contrastive objective and only having nearly 4 million pairs of audio-text data, one epoch is completed in only around 1,000 steps. When training our contrastive model initially, we empirically observe overfitting early in the training process. DALL·E-2 [46] found that using Sharpness-Aware Minimization (SAM) [47] in training CLIP models improves performance; we thus investigate using SAM in our contrastive training stage. SAM is an optimization technique that encourages convergence toward flatter local minima with the hope of improving model generalizability  [47]. SAM is also shown to provide robustness against label noise [47], which is present in our data.
The objective of SAM can be written as:
$\mathcal{L}^{SAM}(\boldsymbol{w})=\max_{\|\boldsymbol{\epsilon}\|_{2}\leq\rho}\mathcal{L}(\boldsymbol{w}+\boldsymbol{\epsilon}),$ (4)
@@
where $\boldsymbol{w}+\boldsymbol{\epsilon}$ is the perturbed weight vector, $\boldsymbol{\epsilon}=\rho\cdot\frac{\nabla_{\boldsymbol{w}}\mathcal{L}(\boldsymbol{w})}{||\nabla_{w}\mathcal{L}(w)||}$ and $\rho$ is a scalar hyperparameter controlling the magnitude of the perturbation. Note that two forward/backward passes are required to compute every update: one for computing the perturbation $\boldsymbol{\epsilon}$ and one for computing the gradient at the perturbed weight vector $\boldsymbol{w}+\boldsymbol{\epsilon}$; thus, the wall-clock time per step is essentially doubled. There has been recent work [48] in reducing SAM’s computational overhead; we leave the investigation of applying these techniques in our setup to future work.
@@
The training dynamics, depicted in Fig. 4, show the impact of incrementally adjusting $\rho$ by increments of 0.025. Increasing $\rho$ corresponds to a larger $\epsilon$ perturbation and thus heavier regularization. The figure clearly shows that without SAM, the model is prone to overfitting on the training data as early as 40,000 steps. In contrast, with even a small degree of sharpness minimization at $\rho=0.025$, overfitting is delayed until approximately 100,000 steps and occurs at a significantly lower objective value. In our final second stage model training, we use SAM with $\rho=0.075$.
@@
Figure 4: Captioning-contrastive objective on the validation set during training, comparing scenarios with and without the application of SAM. For the cases with SAM optimization, we employ various neighborhood sizes as determined by the hyperparameter $\rho$.
@@ V-A Evaluation overview
To evaluate audio-text representations, we focus on three tasks: audio-text retrieval, close-ended audio question answering, and zero-shot transfer in audio classification. These assessments concentrate on the performance of audio-text cross-modal alignment. Additionally, our pretrained audio encoder is also capable of serving as a general-purpose audio representation model for a variety of downstream recognition tasks. To evaluate this, we train a multi-layer perceptron (MLP) on top of our frozen audio encoder to perform audio classification and sound event detection. For open-ended audio understanding tasks that require the generation of free-form text, we evaluate our model on audio captioning tasks. However, we have chosen not to include open-ended audio question answering as it requires an extra text pre-processing and fine-tuning step.
In the audio research community, some major high-impact datasets originate from Freesound, including Clotho [15], ESC50 [49], UrbanSound8K [50], and FSD50K [51]. Another set of popular datasets comes from YouTube, such as Audioset (and its subset AudioCaps) and VGGSound [52]. Given that our dataset includes Freesound and 4 million Youtube samples, we need to ensure that none of the data on which we evaluate is present in our train dataset. FSD50K, a subset of Freesound, is originally in the HEAR benchmark. Because FSD50K is of high-quality and is relatively large, rather than remove FSD50K from our training dataset, we exclude it from our HEAR benchmark evaluation. Mridingham Stroke and Mridingham Tonic also overlaps with Freesound; we choose to exclude it from the HEAR benchmark for simplicity, as it is small-scale relative to other HEAR tasks. Lastly, we exclude Beehive tasks from our HEAR evaluation because they require inference on 600-second long samples, and the models to which we compare are unable to process audio of this length.
V-B Audio-language retrieval
TABLE II: System comparisons on audio-language retrieval on test sets of AudioCaps and Clotho. Higher score means better performance. The evaluation results for the baselines are copied from the references.
@@ V-B3 Result
Our proposed method has achieved state-of-the-art or equivalent performance to the best-performing systems across all evaluated metrics and both datasets. Compared to WavCaps and LAION, our model achieves better performance on AudioCaps and comparable results on Clotho with ‘WavCaps-CNN14’. When compared against FLAP, Cacophony outperforms the non-feature-fusion-based model and matches the fusion-based model, which leverages both global and cropped segments.
TABLE III: Evaluation of audio question answering in Clotho-AQA and Music-AQA benchmarks (“Word” stands for sing-word closed-vocabulary classification, “Bin.” stands for binary classification).
Model Clotho-Word Clotho-Bin. Music-Word Music-Bin.
@@
Ours 19.7 42.6 52.0 70.7 53.6 74.9
V-C Close-ended audio question answering
@@ V-C1 Task definition
Audio question answering (AQA) involves a system taking as input both an audio signal and a natural language question and outputting a response in natural language. We follow the strategy in [28] where the question and answering problem is converted into a supervised classification task, i.e., close-ended audio question answering.

### v3

@@ III Neural Architecture
Figure 3: The proposed system consists of a two-stage training process, as illustrated in the system diagram: Left (Stage 1): Audio-MAE training: This stage is conducted on our collected dataset. After this stage, the audio decoder is discarded, and the encoder is retained to initialize the audio encoder for the second stage of training. Right (Stage 2): Contrastive-captioning training: This stage involves training three components – the audio encoder, the text encoder, and the text decoder. The second stage is dedicated to achieving a contrastive-captioning objective, aligning and fine-tuning the interaction between the audio and text components.
We train Cacophony in two stages, as depicted in Fig. 3. First, we train an audio encoder with the MAE objective, using audio-only data from our collected large-scale dataset. Second, we take this trained audio encoder and use it as the initilzation for training the audio-text model with both contrastive and captioning objectives.
III-A Stage 1: Audio encoder training
We train our audio encoder with an MAE objective, which involves masking random patches of the input signals (e.g., images or spectrograms) and then reconstructing these masked portions [24, 25]. Specifically, in the MAE training, a transformer encoder initially processes the unmasked patches. Following this, a transformer decoder, which will be discarded before the second-stage contrastive-captioning training, predicts both the masked and unmasked regions. This MAE objective encourages the learning of global, contextualized representations over arbitrary subsets of spectrogram patches [25]. Compared to methods based on supervised classification objectives, MAE does not required labeled data. It also demonstrates improved performance with increased model size [24] and increased dataset size [36].
@@ III-A Stage 1: Audio encoder training
III-B Stage 2: Contrastive-captioning training
Given the audio encoder trained in the first stage, we now introduce a text encoder/decoder and train a contrastive model with an auxiliary captioining objective, in a setup similar to BLIP [27] and CoCa [26]. For the contrastive objective [39, 14], we use learned linear projections to map the text and audio embeddings to the same dimension, followed by an $l_{2}$-normalization layer. Then, we use matched audio-text embeddings as positives and all non-paired examples as hard negatives, resulting in $N$ total samples for the pairwise Information Noise-Contrastive Estimation (InfoNCE) loss as described in [40]:
@@ III-B Stage 2: Contrastive-captioning training
where $x_{i}$ and $y_{j}$ are $l_{2}$-normalized embeddings of the audio in the $i$-th pair and text in the $j$-th pair, respectively, and $\tau$ is a learnable temperature.
@@
where $y_{t}$ is the $t$-th ground-truth token for a given caption $y$ and $T$ is the caption’s total length. As a result, we apply both contrastive and generative objectives in the second stage training as follows:
@@
where $\lambda_{\text{Cap}}>0$ is a hyperparameter controlling the relative weight of the captioning loss compared to the contrastive loss.
Architecturally, the second-stage training involves training three key modules: audio encoder, text encoder, and text decoder. For the audio encoder, we use the encoder from the first stage of AudioMAE training, outlined in Sec. III-A. Our text encoder is a transformer that follows RoBERTa’s architecture. However, unlike RoBERTa, we apply causal self-attention rather than bidirectional self-attention. This is a necessary modification to prevent information leakage to the text decoder. We initialize our model with the RoBERTa pretrained weights. The outputs from audio encoder and text encoder are framewise and token-wise embeddings respectively; in order to obtain a single embedding vector for each modality, we choose to integrate multi-head attention poolers on top of the sequential embeddings [41, 26] for the contrastive objective, shown in the right of Fig. 3. Our text decoder consists of stacked transformer layers on top of our text encoder with causal self-attention to prevent next-token prediction leakage. Each transformer block has a cross-attention layer that attends to the full output of our audio encoder. The text decoder layers are initialized randomly.
@@
IV Implementation Details
IV-A Training and inference setup
As outlined in Sec. II, the training datasets are categorized into three types: the ‘clean-labeled’ dataset consists of 1,212 hours, the ‘noisy labeled’ dataset consists of 3,003 hours, and the ‘weakly/unlabeled’ dataset consists of 9,017 hours. All audio files are processed into a mono channel with a sampling rate of 16 kHz. We extract mel-spectrograms using a 25 ms window and a 10 ms hop length, with 128 mel bands and 512 as FFT size. The spectrograms are then converted into 16$\times$16 patches without overlap.
In the first-stage training, we randomly sample 15 seconds of audio for each recording, corresponding to a patch sequence length of 750. At the second stage of training, we limit the audio length to 10.24 seconds, equivalent to a patch sequence length of 512, i.e., when the patch sequence length exceeds 512, we randomly sample 512 patches from the full sequence, following Patchout faSt Spectrogram Transformer (PaSST) [42]. Text samples are tokenized with the pretrained RoBERTa tokenizer and truncated to maximum length of 77. Our audio encoder, employed in both training stages, is a 12-layer ViT-Base (ViT-B) Transformer with 8 attention heads, a hidden size of 768 and an intermediate size of 3072, and utilizes the SiLU activation function. Additionally, a 12-layer decoder with the same settings is used during the MAE training phase. For the attention pooler, we choose to use $n_{head}=8$.
When training AudioMAE, we use a batch size of 512 and a learning rate of $2\times 10^{-4}$, and employ the Adam with decoupled Weight decay (AdamW) [43] optimizer with a weight decay of 0.01. We train for 200,000 steps, with a 10,000 step learning rate warm-up and a cosine decay to $1\times 10^{-6}$. In the contrastive stage, we set $\lambda_{\text{Cap}}=1$ as the default value. While fine-tuning this parameter may marginally improve performance, as suggested by Yu et al. [26], we find the default setting to be sufficiently effective for our purposes. During training, we again use AdamW with weight decay 0.01 with a batch size of 4096. The learning rate is warmed up for 10,000 to a peak of $1\times 10^{-5}$ and is decayed to $1\times 10^{-6}$ over 300,000 (total) steps following a cosine schedule.
During inference, we process the full audio sequence if it fits in the memory. For longer sequences exceeding memory capacity, we randomly sample a subset of spectrogram patches that fit within available memory. This approach allows us to handle variable-length inputs while adapting to computational limitations. Notably, despite the potential for duration distribution shift between training (where we use masked or dropped spectrogram patches) and inference (where we process full sequences), we observe improved performance when utilizing all available audio information during inference, aligning with findings in [23].
@@ IV-B Sharpness-aware minimization
Given our use of a large training batch size of 4096 for the contrastive objective and only having nearly 4 million pairs of audio-text data, one epoch is completed in only around 1,000 steps. When training our contrastive model initially, we empirically observe overfitting early in the training process. DALL·E-2 [44] found that using Sharpness-Aware Minimization (SAM) [45] in training CLIP models improves performance; we thus investigate using SAM in our contrastive training stage. SAM is an optimization technique that encourages convergence toward flatter local minima with the hope of improving model generalizability, which is also shown to provide robustness against label noise [45].
More concretely, we denote $S={(\boldsymbol{x}_{i},y_{i})}^{n}_{i=1}$ as the training dataset and $\boldsymbol{w}$ as the trainable parameters. SAM seeks to find the parameters whose neighbors have low training loss through the following objective:
$\min_{\boldsymbol{w}}\mathcal{L}_{S}^{SAM}(\boldsymbol{w})=\min_{\boldsymbol{w}}\max_{\|\boldsymbol{\epsilon}\|_{p}\leq\rho}\mathcal{L}_{S}(\boldsymbol{w}+\boldsymbol{\epsilon}),$ (4)
where $p\geq 1$ defines the order of the norm and $\rho\geq 0$ is the tunable hyperparameter that measures the size of the neighborhood, and $\mathcal{L}_{S}$ can be any arbitrary loss function on dataset $S$. The inner maximum operator computes the maximum loss within the neighborhood. Its difference from $\mathcal{L}_{S}(\boldsymbol{w})$ defines how quickly the loss increases from $\boldsymbol{w}$, i.e., the sharpness of the loss landscape. The SAM loss is thus minimizing not only the $\mathcal{L}_{S}(\boldsymbol{w})$ itself, but also the sharpness within the neighborhood. To solve the inner maximization problem efficiently, we can use the first-order Taylor expansion approximation, i.e., $\mathcal{L}_{S}(\boldsymbol{w}+\boldsymbol{\epsilon})\approx\mathcal{L}_{S}(\boldsymbol{w})+\boldsymbol{\epsilon}^{T}\nabla_{\boldsymbol{w}}\mathcal{L}_{S}(\boldsymbol{w})$  [45]. With $p=2$, the inner maximization can be achieved at:
$\boldsymbol{\hat{\epsilon}}(\boldsymbol{w})\approx\arg\max_{\|\boldsymbol{\epsilon}\|_{2}\leq\rho}\epsilon^{T}\nabla_{\boldsymbol{w}}\mathcal{L}_{S}(\boldsymbol{w})=\rho\frac{\nabla_{\boldsymbol{w}}\mathcal{L}_{S}(\boldsymbol{w})}{||\nabla_{\boldsymbol{w}}\mathcal{L}_{S}(\boldsymbol{w})||},$ (5)
which is a scaled version of the gradient, normalized to have a magnitude of $\rho$. Substituting $\boldsymbol{\hat{\epsilon}}$ back into equation (4) and differentiating it w.r.t. $\boldsymbol{w}$, we obtain:
$\nabla_{\boldsymbol{w}}\mathcal{L}_{S}^{SAM}(\boldsymbol{w})\approx\nabla_{\boldsymbol{w}}\mathcal{L}_{S}(\boldsymbol{w}+\boldsymbol{\hat{\epsilon}}(\boldsymbol{w}))\approx\nabla_{\boldsymbol{w}}\mathcal{L}_{S}(\boldsymbol{w})|_{(\boldsymbol{w}+\boldsymbol{\hat{\epsilon}})},$ (6)
@@
where the first approximation is due to the Taylor expansion mentioned before, while the second approximation is due the removal of a higher-order differentiation term. Note that two forward/backward passes are required to compute every update: one for computing the perturbation $\hat{\boldsymbol{\epsilon}}$ and the other one for computing the gradient at the perturbed weight vector $\boldsymbol{w}+\hat{\boldsymbol{\epsilon}}$; thus, the wall-clock time per step is essentially doubled. There has been recent work [46] in reducing SAM’s computational overhead; we leave the application of these techniques to future work.
@@
The training dynamics, depicted in Fig. 4, show the impact of incrementally adjusting $\rho$ by increments of 0.025. Increasing $\rho$ corresponds to a larger $\boldsymbol{\epsilon}$ perturbation and thus heavier regularization. The figure clearly shows that without SAM, the model is prone to overfitting on the training data as early as 40,000 steps. In contrast, with even a small degree of sharpness minimization at $\rho=0.025$, overfitting is delayed until approximately 100,000 steps and occurs at a significantly lower objective value. In our final second-stage model training, we use SAM with $\rho=0.075$.
@@
Figure 4: Contrastive-Captioning objective on the validation set during training, comparing scenarios with and without the application of SAM. For the cases with SAM optimization, we employ various neighborhood sizes as determined by the hyperparameter $\rho$.
@@ V-A Evaluation overview
V-A1 Tasks
We evaluate Cacophony’s performance across three primary categories: (1) Audio-Text Representation Tasks: These include audio-text retrieval, closed-ended audio question answering, and zero-shot transfer in audio classification. These assessments concentrate on the model’s performance in audio-text cross-modal alignment. (2) General-Purpose Audio Representation Tasks: To evaluate Cacophony’s pretrained audio encoder as a general-purpose audio representation model, we conduct audio classification and sound event detection tasks. For these assessments, we train a Multi-Layer Perceptron (MLP) on top of our frozen audio encoder. (3) Open-Ended Audio Understanding Task: We assess our model’s capacity of generating free-form text through audio captioning. However, we have excluded open-ended audio question answering from our evaluation, as it requires additional text pre-processing and fine-tuning steps beyond the scope of this study.
V-A2 Evaluation datasets
In the audio research community, some major high-impact datasets originate from Freesound, including Clotho [13], ESC50 [47], UrbanSound8K [48], and FSD50K [49]. Another set of popular datasets comes from YouTube, such as Audioset (and its subset AudioCaps) and VGGSound [50]. Given that our dataset includes Freesound and 4 million Youtube samples, we need to ensure that none of the data on which we evaluate is present in our train dataset. FSD50K, a subset of Freesound, is originally in the HEAR benchmark. Because FSD50K is of high-quality and is relatively large, rather than removing FSD50K from our training dataset, we exclude it from our HEAR benchmark evaluation. Mridingham Stroke and Mridingham Tonic also overlaps with Freesound; we choose to exclude it from the HEAR benchmark for simplicity, as it is small-scale relative to other HEAR tasks. Lastly, we exclude Beehive tasks from our HEAR evaluation because they require inference on 600-second long samples, and the models to which we compare are unable to process audio of this length.
TABLE II: System comparisons on audio-language retrieval on test sets of AudioCaps and Clotho. Results for the baselines are copied from the references. $R@k$ (Recall at $k$) represents the proportion of queries for which the relevant item is retrieved within the top $k$ results. Higher values indicate better performance.
@@ V-B3 Result
Our proposed method has achieved state-of-the-art or comparable performance to the best-performing systems across all evaluated metrics and both datasets. Compared to WavCaps and LAION, our model achieves better performance on AudioCaps and comparable results on Clotho with ‘WavCaps-CNN14’. When compared against FLAP, Cacophony outperforms the non-feature-fusion-based model and matches the fusion-based model. While both Cacophony and FLAP share some common elements, such as LLM-augmented captions and MAE-based audio encoders, our model’s superior performance can be attributed to our two-stage MAE and contrastive training strategy, together with SAM optimization techniques. This combination allows for more effective learning of audio-text representations, resulting in improved retrieval performance across both datasets.
TABLE III: Evaluation of audio question answering in Clotho-AQA and Music-AQA benchmarks (“Word” stands for single-word closed-vocabulary classification, “Bin.” stands for binary classification. “Acc.” stands for accuracy)
[…]
@@
Cacophony (Ours) 19.7 42.6 52.0 70.7 53.6 74.9
V-C Closed-ended audio question answering
@@ V-C1 Task definition
AQA is the task of generating a text response to a text question about an audio signal. We follow the strategy in [29] to cast AQA into a supervised classification task with a closed-ended answer set.